OpenStudy (anonymous):

fin x&y intercepts f(x)=-x^2+3x+10

OpenStudy (bahrom7893):

y - int => set x = 0 y = 0 + 3*0 + 10 = 10, y intercept: (0;10)

OpenStudy (anonymous):

is the x cordinates of the intercept -3,0?

OpenStudy (bahrom7893):

let's check..

OpenStudy (bahrom7893):

of which intercept? x - intercept: y = 0, so: -x^2+3x+10=0

OpenStudy (bahrom7893):

x = -2, x = 5

OpenStudy (anonymous):

x

OpenStudy (anonymous):

when i factored it i used 2 and 5

OpenStudy (bahrom7893):

no, x = -2 and x = 5 Y intercept: (0;10) X intercept: (-2;0) and (5;0)

OpenStudy (anonymous):

ok i see know. thanks for the help.